The number of inmates N (measured in thousands) in federal and state prisons can be approximated by the equation N=-2.02x^2 + 77.89x + 743, where x is the number of years since 1990. In what year was the number of inmates expected to be 1,280,390?
Answer by TimothyLamb(4379)

-2.02x^2 + 77.89x + 743 = 1280390
-2.02x^2 + 77.89x + 1279647 = 0
---
the above quadratic equation is in standard form, with a=-2.02, b=77.89, and c=1279647
---
to solve the above equation using the quadratic formula, plug this:
-2.02 77.89 1279647
into this: https://sooeet.com/math/quadratic-equation-solver.php
---
the roots (x-intercepts) of the above quadratic equation are:
x = 815.433156
x = -776.87375
---
In what year was the number of inmates expected to be 1,280,390?
1990 + 815 = 2805
Answer:
in the year 2805
---

---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php